Find the mass in grams of 2.00 x 10^23 molecules of F2


1
Expert's answer
2022-01-21T15:54:04-0500

Solution:

One mole of any substance contains 6.02×1023 particles.

Thus, one mole of F2 contains 6.02×1023 molecules of F2.

There are 2.00×1023 molecules of F2

Hence,

Moles of F2 = ( 2.00×1023 molecules) × (1 mol / 6.02×1023 molecules) = 0.332 moles


The molar mass of F2 is 38 g/mol

Hence,

Mass of F2 = (0.332 mol) × (38 g / 1 mol) = 12.616 g = 12.62 g


Answer: 12.62 grams of F2